OpenStudy (anonymous):

whatis the pricewise function for f(x) = |2x + 6|?

OpenStudy (anonymous):

\[|x| = \left\{ \begin{matrix}-x, \space x \le 0 \\ x, \space x>0\end{matrix}\right.\]So \(|2x +6| = \left\{ \begin{matrix}-(2x+6), \space 2x+6 \le 0 \\ 2x+6, \space 2x+6>0\end{matrix}\right.\) You are not done. You need to simplify something. Is this clear?

OpenStudy (anonymous):

yes thankyou

OpenStudy (anonymous):

Are you sure you understand completely? I'd glad to clarify more.

OpenStudy (anonymous):

i got it all i have to do is solve the 2x+6 part .

OpenStudy (anonymous):

but o need help with another question :What are the x-intercepts of f(x)= |2x| - 6?

OpenStudy (anonymous):

x-intercepts is when f(x) = 0 so |2x| - 6 = 0. |2x| = 6 Do you know how to get rid of absolute value?